OpenStudy (4meisu):

Solve the equation sin(x+ π/9) = -1/2

OpenStudy (callisto):

sin(x+ π/9) = -1/2 x+ π/9 = 5π /6 or x+ π/9 = 7π / 6 Can you solve them?

OpenStudy (lgbasallote):

uhmm cant you just take the arcsin of both sides?

OpenStudy (anonymous):